OpenStudy (ikou):

Someone please help I'm having trouble understanding this You are going to design an advertisement for a new polynomial identity that you are going to invent. Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships Column A | (x − y), (x + y), (y + x), (y − x)| Column B | (x2 + 2xy + y2), (x2 − 2xy + y2), (ax + b), (cy + d)|

OpenStudy (ikou):

Create your own using the columns below. See what happens when different binomials or trinomials are combined. Square one factor from column A and add it to one factor from column B to develop your own identity.

OpenStudy (aihberkhan):

Are you in FLVS?

OpenStudy (ikou):

yeah this is for algebra II

OpenStudy (aihberkhan):

Step 1: New Identity Create your own using the columns below. See what happens when different binomials or trinomials are combined. Square one factor from column A and add it to one factor from column B to develop your own identity. (ax+b)(x+a) = ax^2 + a^2x + bx + ab Step 2: Proof Check polynomial identity. (ax+b)(x+a) = FOIL (ax+b)(x+a) ax^2+a^2x is the First Term in the FOIL ax^2 + a^2x + bx + ab (ax+b)(x+a) +bx+ab is the Second Term in the FOIL Then add both expressions together from First Term and Second Term = ax^2 + a^2x + bx + ab Step 3: Proof Try with numbers now (ax+b)(x+a) = ax^2 + a^2x + bx + ab ((2*5)+8)(5+2) =(2*5^2)+(2^2*5)+(8*5)+(2*8) ((10)+8)(7) =(2*25)+(4*5)+(40)+(16) (18)(7) =(50)+(20)+(56) 126 =126

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @Ikou

OpenStudy (ikou):

(x+a) isn't in any of the columns though